OpenStudy (anonymous):

How would I evaluate the following limit:

OpenStudy (anonymous):

\[\Huge Evaluate:\lim_{x \rightarrow 1} \frac {x^3-1}{x^2-1}?\]

OpenStudy (unklerhaukus):

\[\lim_{x \rightarrow 1} \frac {x^3-1}{x^2-1}\]\[ ^{L'H}=\lim_{x \rightarrow 1} \frac {3x^2}{2x}\]\[ ^{L'H}=\lim_{x \rightarrow 1} \frac {6x}{2}\]\[=\lim_{x\rightarrow 1}3x\]\[=\]

OpenStudy (anonymous):

Where does the three and two come from?

OpenStudy (accessdenied):

I would try factoring the numerator and the denominator using these little formulas and then cancelling common expression in the numerator / denominator: a^2 - b^2 = (a + b)(a - b) a^3 - b^3 = (a - b)(a^2 + ab + b^2

OpenStudy (anonymous):

\[(x^3 - b^3)= (x-b)(x^2 + bx + b^2)\] \[(x^2 - b^2)= (x-b)(x+b)\] So how about you factor the top and bottom

OpenStudy (unklerhaukus):

L'Hôpital's rule when the limit of a fraction has infinities or zeros you can differentiate both the numerator and denominator and the limit of this will be equal to the true measure of the limit

OpenStudy (anonymous):

I'm in Honors PreCalculus...

OpenStudy (anonymous):

Ok then just factor as we told you if you haven't learnt L'Hôpital's rule

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}((x-1)(x^2 +x+1^2))/((x+1)(x-1))\]

OpenStudy (anonymous):

Can you go on now?

OpenStudy (anonymous):

Yes, it makes sense now.

OpenStudy (anonymous):

Great.

OpenStudy (unklerhaukus):

L'Hôpital's rule can be applied if \[\lim_{x\rightarrow c}={f(x)}=\lim_{x\rightarrow c}={g(x)}=0 \text{ or } \pm \infty\] \[\lim_{x\rightarrow c}=\frac{f(x)}{g(x)}\qquad^{L'H}=\lim_{x\rightarrow c}=\frac{f'(x)}{g'(x)}\]

OpenStudy (anonymous):

Thank You, @UnkleRhaukus & @AccessDenied !

OpenStudy (anonymous):

And @beeqay !

OpenStudy (accessdenied):

I'm not sure you could use L'Hopital's rule that second time since x->1 of that function was already in form 3x^2 / x, and x->1 didn't result in any issues. You're welcome. :)

OpenStudy (unklerhaukus):

typo** \[\lim_{x\rightarrow c}{f(x)}=\lim_{x\rightarrow c}{g(x)}=0 \text{ or } \pm \infty\]

OpenStudy (anonymous):

No problem.